OpenStudy (anonymous):

Evaluate P(3, 0). 3!/0!(3-0)! is it 1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

But your formula is wrong..

OpenStudy (anonymous):

0! = 1

OpenStudy (anonymous):

yep:P sorry

OpenStudy (anonymous):

\[\huge ^3P_0 = \frac{3!}{(3-0)!} \implies \frac{3!}{3!} = 1\]